The active mass of 7g of nitrogen ina 2l container wo uld be ?

Grade:11

2 Answers

Arun
4 years ago

 

Moles of N2 = 7/28 = 0.25 mol
Active mass = conc. in mol L-1
                  = 0.25 mol / 2 L
                  = 0.125
 
In case of any query please feel free to ask
 
Regards
Arun
Vikas TU
4 years ago
Active mass is the concentration of the compound per unit volume in liter or active mass can be determined exactly in the way to calculate concentration per unit volume of solute.
Its unit is mol /L
n = m /M
n = 7 /28
= 0.25 mol
where, n is Moles of N2
m is mass of N2
M is molar mass
Active mass = conc. in mol L-1
                  = 0.25 mol / 2 L
                  = 0.125